OpenStudy (anonymous):

The temperature on a particular day started at –13°F. It rose steadily by 3° each hour. The function y = –13 + 3x models the temperature, where x is the number of hours and y is the temperature. Graph the equation.

OpenStudy (anonymous):

just plug in numbers for x and you will get it

OpenStudy (anonymous):

wait a minute..,

OpenStudy (anonymous):

i am using matlab now

OpenStudy (anonymous):

sorry my computer froze. im back though

OpenStudy (anonymous):

so i just put in any numbers?

OpenStudy (anonymous):

yes it should be linear y = –13 + 3x plug x=0;y = -13+3*0=-13 the graph i put shows -10 so you can get the rest

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

you should get something like this